# SphereDrag: Spherical Geometry-Aware Panoramic Image Editing

### v1

@@ Abstract
Image editing has made great progress on planar images, but panoramic image editing remains underexplored. Due to their spherical geometry and projection distortions, panoramic images present three key challenges: boundary discontinuity, trajectory deformation, and uneven pixel density. To tackle these issues, we propose SphereDrag, a novel panoramic editing framework utilizing spherical geometry knowledge for accurate and controllable editing. Specifically, adaptive reprojection (AR) uses adaptive spherical rotation to deal with discontinuity; great-circle trajectory adjustment (GCTA) tracks the movement trajectory more accurate; spherical search region tracking (SSRT) adaptively scales the search range based on spherical location to address uneven pixel density. Also, we construct PanoBench, a panoramic editing benchmark, including complex editing tasks involving multiple objects and diverse styles, which provides a standardized evaluation framework. Experiments show that SphereDrag gains a considerable improvement compared with existing methods in geometric consistency and image quality, achieving up to 10.5% relative improvement.
keywords:
Interactive Point-based Editing , Diffusion Models , Panoramic Image
journal: Neurocomputing
[…]
@@ 1 Introduction
Image generation is one of the current research hotspots. Many works [1, 2, 3, 4, 5] have made significant progress in controllable high-quality planar image generation, primarily by fine-tuning pre-trained large-scale diffusion models to fit various application scenarios. Numerous studies [6, 7, 8] have also been proposed for image editing, enabling highly controllable modifications to existing images based on new requirements. To gain a more comprehensive understanding of the scene, spherical panoramic images, also known as omnidirectional panoramic images or 360 panoramic images, are widely used in autonomous driving [9, 10, 11], and virtual reality [12], etc.
@@ 2.1 Image Editing
Image generation is a popular topic in image generation because of its user-friendly interaction. Early image editing methods are based on Generative Adversarial Networks (GANs) [13, 14, 15, 16, 17], which made great progress in image generation but struggled with training instability and mode collapse [18, 19, 20]. These limitations make researchers explore more advanced generation models. Diffusion models [21, 22] become popular because of its ability to generate high-quality and highly controllable images through iterative denoising, and achieve remarkable improvements in both fidelity and diversity [23, 24, 25, 26].
Recently, many works focus on sophisticated image editing capabilities [27, 28, 29]. Early ones integrate natural language with generation models to enable text-guided editing [30, 31, 32], but these methods suffer from insufficient precision due to inherent textual ambiguity [1], limiting their effectiveness for fine-grained adjustments. To deal with these challenges, researchers develop more precise point-driven editing methods based on interactive drag operations. DragGAN [33] pioneered this approach by enabling intuitive local structure adjustments with direct user control. Building upon this foundation, subsequent works incorporate diffusion framework [34], leading to DragDiffusion [6], which leverages diffusion models’ advantages in maintaining global consistency while enabling local edits. DragNoise [8] further refine this paradigm through innovative noise optimization strategies that better balance local control with global coherence. StableDrag [7] introduces a robust framework for point-based image editing by enhancing DragGAN and DragDiffusion with a discriminative point tracking method and a confidence-based latent enhancement strategy.
@@ 2.2 Spherical Panoramic Image Generation
In the domain of GAN-based [13, 14, 15, 16, 17] approaches, COCO-GAN [35] introduces a coordinate-conditional framework that generates images in a divided manner, using spatial coordinates as a guiding signal for the generator to progressively synthesize local regions. Its discriminator evaluates global consistency, local appearance, and edge continuity, ensuring seamless output and enabling prediction beyond training dimensions via cylindrical coordinates. PanoGAN [36] utilizes an adversarial feedback mechanism along with a cross-view generation framework, achieving a breakthrough in the semantic matching between aerial perspective and ground-level panoramic images. By using a dual-branch discriminator and repeated optimization, it can ensure the feature matching at the pixel level and maintain the spatial consistency within multimodal data. Similarly, SphericGAN [37] adopts a semi-supervised hyperspherical latent space modeling strategy, aligning real and generated data clusters to capture category-independent variations effectively . Furthermore, CycleGAN [38, 39] uses a cycle-consistent adversarial network, and addresses unpaired training data scenarios, making it possible to perform cross-domain image translation without explicit pixel-wise correspondences.
Table 1: Notations
$sg(\cdot)$ Stop gradient, preventing gradient backpropagation
$P_{han}/P_{tar}$ The handle point / The target point
$P_{mid}$ The midpoint between the handle and target point
$P_{han}^{k}$ The current handle point at the $k$-th iteration
$r$ The base radius determining the search region size
$(\theta,\phi)$ The spherical coordinates with longitude $\theta$ and latitude $\phi$
$\mathcal{F}$ The feature maps
$\Omega_{search}$ The search region
$\hat{w}_{t}^{k}$ The latent variable at fixed timestep $t$ after $k$-th iteration
$w^{0}_{t}$ The initial latent variable at timestep $t$
$T_{\text{Sphere}\rightarrow\text{2D}}(\cdot)$ The projection from spherical coordinates $(\theta,\phi)$ to 2D
$T_{\text{2D}\rightarrow\text{Sphere}}(\cdot)$ The projection from 2D to spherical coordinates $(\theta,\phi)$
$M$ A binary mask indicating editable / uneditable regions
$\vec{d}$ The direction from the current handle point to the target point
$\mathcal{L}_{\text{ms}}$ The motion supervision loss
$F_{q}(\cdot)$ The feature maps extracted at location $q$
$F_{q}(\hat{w}_{t}^{k})$ The feature maps extracted from $\hat{w}_{t}^{k}$ at $q$ with spherical
alignment and resizing
$\mathcal{R}_{\text{align}}(\theta,\phi)$ The rotation to align $(\theta,\phi)$ with a suitable panoramic representation
$\mathcal{G}_{\text{gc}}(\cdot)$ The great-circle movement direction computation
@@ 3 Method
In this section, we introduce our SphereDrag in detail. First, we briefly introduce classic image editing algorithm and build up our notations. The notations are summarized in table 1. Next, we introduce our SphereDrag framework. Faced with boundary discontinuity challenge, we introduce our adaptive reprojection (AR) to get the suitable panoramic representation. Also, we introduce our spherical latent optimization, including great-circle trajectory adjustment (GCTA) and spherical search region tracking (SSRT), to deal with trajectory deformation and uneven pixel density, respectively.
@@ 3.1 Background and Notations
$\mathcal{L}_{\text{ms}}(\hat{w}_{t}^{k})=$ $\sum_{q\in\mathcal{F}}\left\|F_{q+\vec{d}}(\hat{w}_{t}^{k})-\text{sg}(F_{q}(\hat{w}_{t}^{k}))\right\|_{1}$
@@
where $\mathcal{L}_{\text{ms}}$ denotes the motion supervision loss and guides how the latent code $\hat{w}_{t}^{k}$ is updated during movement. The first term encourages the feature maps to move from its current position to a target position specified by the motion direction $\vec{d}$, $F_{q}(\cdot)$ denotes the feature extracted at spatial location $q$ from the latent code, $q\in\mathcal{F}$ refers to a spatial coordinate $(x,y)$ within the region $\mathcal{F}$. $P_{han}^{k}$ denotes the handle point at the $k$-th iteration. The second term constrains the uneditable region to remain unchanged, where $M$ denotes a binary mask with $M=1$ indicating editable regions and $M=0$ indicating uneditable regions, $\hat{w}_{t-1}^{k}$ denotes the latent code at fixed timestep $t-1$ after $k$-th iteration, and $\hat{w}_{t-1}^{0}$ denotes its initial version before optimization. It penalizes the difference between the current latent $\hat{w}_{t-1}^{k}$ and its initial version $\hat{w}_{t-1}^{0}$, helping the uneditable region stay still. $\lambda$ controls the strength of the constraint that keeps the uneditable region unchanged, and $\odot$ denoting the element-wise multiplication operation. $\text{sg}(\cdot)$ denotes the stop-gradient operation, which treats the feature maps $F_{q}(\hat{w}_{t}^{k})$ and the initial latent code $\hat{w}_{t-1}^{0}$as fixed targets during optimization, preventing their gradients from affecting the update of $\hat{w}_{t}^{k}$.
@@ 3.2.1 Adaptive Reprojection (AR)
where the function $\mathcal{T}_{\text{align}}(\cdot)$ denotes the complete alignment pipeline. $I^{\prime}$ denotes the image, $M^{\prime}$ denotes the mask, and $\mathcal{P}^{\prime}$ denotes the handle point and target point. $T_{\text{2D}\rightarrow\text{Sphere}}(\cdot)$ is the transformation that projects the input data $(I,M,\mathcal{P})$ from 2D coordinates to spherical coordinates. The rotation matrix $R$ is applied to these spherical coordinates to align them with a suitable panoramic representation. Finally, $T_{\text{Sphere}\rightarrow\text{2D}}(\cdot)$ projects the rotated data back to 2D coordinates, resulting in the final image, mask,handle point and target point, $(I^{\prime},M^{\prime},\mathcal{P}^{\prime})$. As shown in fig. 3, our spherical alignment strategy makes the editing region more stable semantic consistency and trajectory continuity. The further details are provided in appendix A.
@@ 3.2.2 Great-Circle Trajectory Adjustment (GCTA)
where $\mathcal{G}_{\text{gc}}(\cdot)$ denotes the great-circle direction computation function (see  appendix B for details).
@@
$\mathcal{L}_{\text{ms}}(\hat{w}_{t}^{k})=$ $\sum_{q\in\mathcal{F}}\left\|F_{q+\vec{d}}(\hat{w}_{t}^{k})-\text{sg}(F_{q}(\hat{w}_{t}^{k}))\right\|_{1}$
@@
Figure 4: In a panoramic image, $\cos\phi$ acts as the stretching factor, ensuring that the degree of stretching in the latitude direction from the sphere to the plane is consistent with the spherical geometry. The variation of $\cos\phi$ with latitude $\phi$ precisely reflects the projection requirements of the parallel lengths on the sphere, thus maintaining the accuracy and consistency of geographic information during the projection process.
@@ 3.2.3 Spherical Search Region Tracking (SSRT)
where $r$ is a fixed base radius. The use of $\cos\phi$ compensates for the horizontal stretching caused by the ERP in high-latitude areas, improving the geometric consistency of the search region and improving the accuracy of the tracking. An illustration of how $\cos\phi$ relates to latitude distortion is provided in fig. 4.
@@ 4 Experiments
In this section, we conduct experiments to evaluate SphereDrag. Firstly, we introduce the benchmark, experimental protocols, and evaluation metrics used in our experiments. Secondly, we compare the performance of SphereDrag with state-of-the-art image editing methods. Thirdly, we carry out ablation experiments to clarify the contribution of each module. Finally, we provide a comprehensive analysis and discussion of the results.
Table 2: Quantitative results of our method (Ours) and state-of-the-art models, Baseline (DragDiffusion), DragNoise, StableDrag. They are evaluated on images with varying FOV: 30, 60, and 90. Metrics include IF (higher is better), FID, and sFID (lower is better).
[…]
Baseline 30 0.6666 125.6694 409.8609
DragNoise 0.7161 126.3004 391.0360
StableDrag 0.7133 116.4910 390.5301
Ours 0.7364 114.8911 375.9909
Baseline 60 0.8037 80.4527 277.0527
DragNoise 0.8047 88.8964 302.2573
StableDrag 0.8312 72.2143 253.5878
Ours 0.8495 70.4464 244.3402
Baseline 90 0.8924 64.0885 182.5420
DragNoise 0.8734 77.1288 226.8186
StableDrag 0.9064 57.7487 168.7382
Ours 0.9191 52.8041 160.2049
@@ 4.1 Benchmark, Protocols, and Evaluation Metrics
Evaluation Metrics Image Fidelity (IF), Fréchet Inception Distance (FID), and Spatial Fréchet Inception Distance (sFID) are used as evaluation metrics.
@@
  • 1.
[…]
  • 2.
    $\text{FID}=\|\mu_{1}-\mu_{2}\|_{2}^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{1/2}\right),$ (13)
  • 3.
Table 3: Ablation study results comparing the performance of different model variants, evaluated on images with varying FOV of 30, 60, and 90. AR / GCTA / SSRT indicates our adaptive reprojection / great-circle trajectory adjustment / spherical search region tracking, respectively.
[…]
30 0.6666
0.7060
0.7081
0.7364
60 0.8037
0.8321
0.8326
0.8495
90 0.8924
0.9101
0.9103
0.9191
Experimental Protocols All experiments are based on Stable Diffusion v1.5 with a uniform learning rate of 0.01. We set the latent variable at fixed timestep $t$ = 35, the strength of the constraint $\lambda$ = 0.1, the suitable panoramic representation, we set the longitude to 0, The latitude remains consistent before and after rotation and maintain a resolution of the $1024\times 512$ image. The implementation runs on servers equipped with 8 NVIDIA L40S GPUs. The rest of the protocols are the same as DragDiffusion.
Figure 6: Ablation study on adaptive reprojection (AR). The top row shows the result with AR enabled, while the bottom row shows the result without it. The red boxes highlight the seam regions. With AR, the transitions across the seams are smooth and visually coherent. In contrast, without AR, noticeable artifacts and discontinuities appear. These results demonstrate the effectiveness of AR in enhancing geometric alignment and visual consistency in panoramic image editing.

### v2

@@ Abstract
Image editing has made great progress on planar images, but panoramic image editing remains underexplored. Due to their spherical geometry and projection distortions, panoramic images present three key challenges: boundary discontinuity, trajectory deformation and uneven pixel density. To tackle these issues, we propose SphereDrag, a novel panoramic editing framework utilizing spherical geometry knowledge for accurate and controllable editing. Specifically, adaptive reprojection (AR) uses adaptive spherical rotation to deal with discontinuity; great-circle trajectory adjustment (GCTA) tracks the movement trajectory more accurate; spherical search region tracking (SSRT) adaptively scales the search range based on spherical location to address uneven pixel density. Also, we construct PanoBench, a panoramic editing benchmark, including complex editing tasks involving multiple objects and diverse styles, which provides a standardized evaluation framework. Experiments show that SphereDrag gains a considerable improvement compared with existing methods in geometric consistency and image quality, achieving up to 10.5% relative improvement.
@@ 1 Introduction
Image generation is one of the current research hotspots. Many works [18, 27, 29, 9] have made significant progress in controllable high-quality planar image generation, primarily by fine-tuning pre-trained large-scale diffusion models to fit various application scenarios. Numerous studies [24, 4, 16] have also been proposed for image editing, enabling highly controllable modifications to existing images based on new requirements. To gain a more comprehensive understanding of the scene, spherical panoramic images, also known as omnidirectional panoramic images or 360 panoramic images, are widely used in autonomous driving [13, 14], and virtual reality, etc.
@@ 2.1 Image Editing
Image generation is a popular topic in image generation because of its user-friendly interaction. Early image editing methods are based on Generative Adversarial Networks (GANs) [6, 11], which made great progress in image generation but struggled with training instability and mode collapse [17, 1]. These limitations make researchers explore more advanced generation models. Diffusion models [8] become popular because of its ability to generate high-quality and highly controllable images through iterative denoising, and achieve remarkable improvements in both fidelity and diversity [23, 22].
Recently, many works focus on sophisticated image editing capabilities [12, 2]. Early ones integrate natural language with generation models to enable text-guided editing [21, 5], but these methods suffer from insufficient precision due to inherent textual ambiguity [27], limiting their effectiveness for fine-grained adjustments. To deal with these challenges, researchers develop more precise point-driven editing methods based on interactive drag operations. DragGAN [20] pioneered this approach by enabling intuitive local structure adjustments with direct user control. Building upon this foundation, subsequent works incorporate diffusion framework [30], leading to DragDiffusion [24], which leverages diffusion models’ advantages in maintaining global consistency while enabling local edits. DragNoise [16] further refine this paradigm through innovative noise optimization strategies that better balance local control with global coherence. StableDrag [4] introduces a robust framework for point-based image editing by enhancing DragGAN and DragDiffusion with a discriminative point tracking method and a confidence-based latent enhancement strategy.
@@ 2.2 Spherical Panoramic Image Generation
In the domain of GAN-based [6, 11] approaches, COCO-GAN [15] introduces a coordinate-conditional framework that generates images in a divided manner, using spatial coordinates as a guiding signal for the generator to progressively synthesize local regions. Its discriminator evaluates global consistency, local appearance, and edge continuity, ensuring seamless output and enabling prediction beyond training dimensions via cylindrical coordinates. PanoGAN [25] utilizes an adversarial feedback mechanism along with a cross-view generation framework, achieving a breakthrough in the semantic matching between aerial perspective and ground-level panoramic images. By using a dual-branch discriminator and repeated optimization, it can ensure the feature matching at the pixel level and maintain the spatial consistency within multimodal data. Similarly, SphericGAN [3] adopts a semi-supervised hyperspherical latent space modeling strategy, aligning real and generated data clusters to capture category-independent variations effectively. Furthermore, CycleGAN [32] uses a cycle-consistent adversarial network, and addresses unpaired training data scenarios, making it possible to perform cross-domain image translation without explicit pixel-wise correspondences.
@@ 3 Method
In this section, we introduce our SphereDrag in detail. First, we briefly introduce classic image editing algorithm and build up our notations. The notations are summarized in Section A “Notations used in our method” in the Supplementary Material.
Next, we introduce our SphereDrag framework. Faced with boundary discontinuity challenge, we introduce our adaptive reprojection (AR) to get the suitable panoramic representation. Also, we introduce our spherical latent optimization, including great-circle trajectory adjustment (GCTA) and spherical search region tracking (SSRT), to deal with trajectory deformation and uneven pixel density, respectively.
@@ 3.1 Background and Notations
$\mathcal{L}_{\text{ms}}(\hat{w}_{t}^{k})=\$$\sum_{q\in\mathcal{F}}\left\|F_{q+\vec{d}}(\hat{w}_{t}^{k})-\text{sg}(F_{q}(\hat{w}_{t}^{k}))\right\|_{1}$
@@
The motion supervision loss $\mathcal{L}_{\text{ms}}$ guides the update of the latent code $\hat{w}_{t}^{k}$ during movement. The first term encourages feature maps extracted at spatial locations $q\in\mathcal{F}$ to move toward a target position along the motion direction $\vec{d}$, where $F_{q}(\cdot)$ denotes the feature at location $q=(x,y)$, and $P_{\text{han}}^{k}$ is the handle point at the $k$-th iteration. The second term enforces consistency in uneditable regions. A binary mask $M$ indicates editable ($M=1$) and uneditable ($M=0$) areas. It penalizes deviations between the current latent code $\hat{w}_{t-1}^{k}$ and its initial version $\hat{w}_{t-1}^{0}$ only in the uneditable regions ($M=0$), with the strength controlled by $\lambda$ and using element-wise multiplication $\odot$. The stop-gradient operator $\text{sg}(\cdot)$ prevents gradients from backpropagating through fixed reference features $F_{q}(\hat{w}_{t}^{k})$ and $\hat{w}_{t-1}^{0}$.
@@ 3.2.1 Adaptive Reprojection (AR)
where the function $\mathcal{T}_{\text{align}}(\cdot)$ denotes the complete alignment pipeline. $I^{\prime}$ denotes the image, $M^{\prime}$ denotes the mask, and $\mathcal{P}^{\prime}$ denotes the handle point and target point. $T_{\text{2D}\rightarrow\text{Sphere}}(\cdot)$ is the transformation that projects the input data $(I,M,\mathcal{P})$ from 2D coordinates to spherical coordinates. The rotation matrix $R$ is applied to these spherical coordinates to align them with a suitable panoramic representation. Finally, $T_{\text{Sphere}\rightarrow\text{2D}}(\cdot)$ projects the rotated data back to 2D coordinates, resulting in the final image, mask,handle point and target point, $(I^{\prime},M^{\prime},\mathcal{P}^{\prime})$. As shown in fig. 3, our spherical alignment strategy makes the editing region more stable semantic consistency and trajectory continuity. The further details are provided in Section B “Detail for $\mathcal{T}_{\text{align}}(\cdot)$ in Adaptive Reprojection (AR)” in the Supplementary Material.
@@ 3.2.2 Great-Circle Trajectory Adjustment (GCTA)
where $\mathcal{G}_{\text{gc}}(\cdot)$ denotes the great-circle direction computation function (see Section C “Detail for Great Circle Direction Calculation” in the Supplementary Material).
@@
$\mathcal{L}_{\text{ms}}(\hat{w}_{t}^{k})=\$$\sum_{q\in\mathcal{F}}\left\|F_{q+\vec{d}}(\hat{w}_{t}^{k})-\text{sg}(F_{q}(\hat{w}_{t}^{k}))\right\|_{1}$
@@ 3.2.3 Spherical Search Region Tracking (SSRT)
where $r$ is a fixed base radius. The use of $\cos\phi$ compensates for the horizontal stretching caused by the ERP in high-latitude areas, improving the geometric consistency of the search region and improving the accuracy of the tracking. An illustration of how $\cos\phi$ relates to latitude distortion is provided in Section D “Stretching Factor and Latitude Distortion” in the Supplementary Material.
@@ 4 Experiments
In this section, we conduct experiments to evaluate SphereDrag. Firstly, we introduce the benchmark, experimental protocols, and evaluation metrics used in our experiments. Secondly, we compare the performance of SphereDrag with state-of-the-art image editing methods. Third, we carry out ablation experiments to clarify the contribution of each module. Finally, we provide a detailed discussion of the experimental hyperparameters in Section E “Discussion” in the Supplementary Material.
@@ 4.1 Benchmark, Protocols, and Evaluation Metrics
Evaluation Metrics Image Fidelity (IF), Fréchet Inception Distance (FID), and Spatial Fréchet Inception Distance (sFID) are used as evaluation metrics. Detailed definitions and formulations of these metrics are provided in Section F “Evaluation Metrics” in the Supplementary Material.
Experimental Protocols All experiments are based on Stable Diffusion v1.5 with a uniform learning rate of 0.01. To enhance the model’s fidelity for panoramic scenes, the training process involves fine-tuning the cross-attention layers of its UNet architecture via Low-Rank Adaptation (LoRA). We set the latent variable at fixed timestep $t$ = 35, the strength of the constraint $\lambda$ = 0.1, the suitable panoramic representation, we set the longitude to 0, The latitude remains consistent before and after rotation and maintain a resolution of the $1024\times 512$ image. The implementation runs on servers equipped with 8 NVIDIA L40S GPUs. The rest of the protocols are the same as DragDiffusion.
@@ 4.2 Performance Comparison
Table 1: Quantitative results of our method (Ours) and state-of-the-art models. All methods are evaluated on images with FOV of 30, 60, and 90. Metrics include IF (higher is better), FID, and sFID (lower is better).
Method 30 60 90
[…]
Baseline 0.6666 125.6694 409.8609 0.8037 80.4527 277.0527 0.8924 64.0885 182.5420
DragNoise 0.7161 126.3004 391.0360 0.8047 88.8964 302.2573 0.8734 77.1288 226.8186
StableDrag 0.7133 116.4910 390.5301 0.8312 72.2143 253.5878 0.9064 57.7487 168.7382
Ours 0.7364 114.8911 375.9909 0.8495 70.4464 244.3402 0.9191 52.8041 160.2049
4.3 Ablation Studies
In this section, we analyze the contribution of each module to overall performance on images with FOV angles of 30, 60, and 90. The evaluation is conducted using the IF metric. To provide a more thorough assessment, we also incorporate the sFID metric in an extended experiment.
The results show that AR consistently improves performance across all FOV settings. Adding GCTA brings further gains, and combining AR, GCTA, and SSRT achieves the best results. Section G, “Effectiveness of Adaptive Reprojection,” in the Supplementary Material provides visualizations demonstrating that AR significantly enhances semantic continuity and geometric structure in the output panoramic images.
Table 2: Ablation study results comparing different model variants on images with FOV of 30, 60, and 90. AR / GCTA / SSRT denote adaptive reprojection, great-circle trajectory adjustment, and spherical search region tracking, respectively.
[…]
30 60 90 30 60 90
0.6666 0.8037 0.8924 409.86 277.05 182.54
0.7060 0.8321 0.9101 398.08 260.21 172.47
0.7081 0.8326 0.9103 399.30 259.97 172.21
0.7364 0.8495 0.9191 375.99 244.34 160.20
5 Conclusion
Acknowledgements
This work is supported in part by Natural Science Foundation of Shanghai under Grant 24ZR1425600, “Pioneer” and “Leading Goose” R&D Program of Zhejiang (No.2025C02014), Ningbo Science and Technology Special Projects under Grant No. 2025Z028, the Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission with Grant No. 24CGA73, and the Fundamental Research Funds for the Central Universities.
Supplementary Material: “SphereDrag: Spherical Geometry-Aware Panoramic Image Editing”
A Notations used in our method
$sg(\cdot)$ Stop gradient operation, preventing gradient backpropagation.
$P_{\text{han}}/P_{\text{tar}}$ The handle point / the target point.
$P_{\text{mid}}$ The midpoint between the handle and target point.
$P_{\text{han}}^{k}$ The current handle point at the $k$-th iteration.
$r$ The base radius determining the search region size.
$(\theta,\phi)$ The spherical coordinates: longitude $\theta$, latitude $\phi$.
$\mathcal{F}$ The feature maps.
$\Omega_{\text{search}}$ The search region.
$\hat{w}_{t}^{k}$ The latent variable at fixed timestep $t$ after the $k$-th iteration.
$w_{t}^{0}$ The initial latent variable at timestep $t$.
$T_{\text{Sphere}\rightarrow\text{2D}}(\cdot)$ The projection from spherical to 2D coordinates.
$T_{\text{2D}\rightarrow\text{Sphere}}(\cdot)$ The projection from 2D to spherical coordinates.
$M$ A binary mask indicating editable vs. uneditable regions.
$\vec{d}$ The direction vector from the current handle point to the target point.
$\mathcal{L}_{\text{ms}}$ The motion supervision loss.
$F_{q}(\cdot)$ Feature maps extracted at location $q$.
$F_{q}(\hat{w}_{t}^{k})$ Feature maps extracted from $\hat{w}_{t}^{k}$ at $q$, with spherical alignment and resizing.
$\mathcal{R}_{\text{align}}(\theta,\phi)$ Rotation to align $(\theta,\phi)$ with panoramic view.
$\mathcal{G}_{\text{gc}}(\cdot)$ Great-circle direction computation.
B Detail for $\mathcal{T}_{\text{align}}(\cdot)$ in Adaptive Reprojection (AR)
$\theta=\frac{\pi}{2}-\frac{j}{H}\cdot\pi,\quad\phi=\frac{i}{W}\cdot 2\pi-\pi.$ (S-1)
$R_{\phi}(\Delta\phi)=\begin{bmatrix}\cos(\Delta\phi)&-\sin(\Delta\phi)&0\\ \sin(\Delta\phi)&\cos(\Delta\phi)&0\\ 0&0&1\end{bmatrix}$, (S-2)
$R_{\theta}(\Delta\theta)=\begin{bmatrix}1&0&0\\ 0&\cos(\Delta\theta)&-\sin(\Delta\theta)\\ 0&\sin(\Delta\theta)&\cos(\Delta\theta)\end{bmatrix}$, (S-3)
$R(\Delta\theta,\Delta\phi)=R_{\theta}(\Delta\theta)\cdot R_{\phi}(\Delta\phi),$ (S-4)
C Detail for Great Circle Direction Calculation
$\theta=\frac{\pi}{2}-\frac{j}{H}\cdot\pi,\quad\phi=\frac{i}{W}\cdot 2\pi-\pi,$ (S-5)
$x$ $=\cos(\theta)\cos(\phi),$ (S-6)
$y$ $=\cos(\theta)\sin(\phi),$
$z$ $=\sin(\theta).$
$\mathbf{n}=\frac{\mathbf{P}_{\text{han}}\times\mathbf{P}_{\text{tar}}}{\|\mathbf{P}_{\text{han}}\times\mathbf{P}_{\text{tar}}\|}.$ (S-7)
$\mathbf{P}^{\prime}=\frac{\mathbf{P}_{\text{han}}^{k}-(\mathbf{P}_{\text{han}}^{k}\cdot\mathbf{n})\mathbf{n}}{\|\mathbf{P}_{\text{han}}^{k}-(\mathbf{P}_{\text{han}}^{k}\cdot\mathbf{n})\mathbf{n}\|}.$ (S-8)
$\mathbf{v}_{\text{move}}=\frac{\mathbf{P}_{\text{tar}}-\mathbf{P}^{\prime}}{\|\mathbf{P}_{\text{tar}}-\mathbf{P}^{\prime}\|}.$ (S-9)
$\hat{\mathbf{e}}_{\phi}$ $=\frac{1}{\cos(\theta)}\frac{\partial\mathbf{P}}{\partial\phi}=\left[-\sin(\phi),\ \cos(\phi),\ 0\right]^{T},$ (S-10)
$\hat{\mathbf{e}}_{\theta}$ $=\frac{\partial\mathbf{P}}{\partial\theta}=\left[-\sin(\theta)\cos(\phi),\ -\sin(\theta)\sin(\phi),\ \cos(\theta)\right]^{T}.$ (S-11)
$\begin{bmatrix}v_{\phi}\\ v_{\theta}\par\end{bmatrix}=\begin{bmatrix}\mathbf{v}_{\text{move}}\cdot\hat{\mathbf{e}}_{\phi}\\ \mathbf{v}_{\text{move}}\cdot\hat{\mathbf{e}}_{\theta}\par\end{bmatrix}$. (S-12)
$\vec{d}=\frac{[\frac{v_{\phi}}{\cos(\theta)},v_{\theta}]}{\left\|\left[\frac{v_{\phi}}{\cos(\theta)},v_{\theta}\right]\right\|}.$ (S-13)
D Stretching Factor and Latitude Distortion
Figure S-1: In a panoramic image, $\cos\phi$ acts as the stretching factor, ensuring that the degree of stretching in the latitude direction from the sphere to the plane is consistent with the spherical geometry. The variation of $\cos\phi$ with latitude $\phi$ precisely reflects the projection requirements of the parallel lengths on the sphere, thus maintaining the accuracy and consistency of geographic information during the projection process.
E Evaluation Metrics
    $\text{IF}=1-\text{avg (LPIPS)},$ (S-14)
    $\text{FID}=\|\mu_{1}-\mu_{2}\|_{2}^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{1/2}\right),$ (S-15)
Figure S-2: PCA visualization results. At smaller timesteps (e.g., $t=1,10$), the latent variable contains contains redundant details, while at larger timesteps (e.g., $t=50$), the spatial information is overly diminished. At $t=35$, a good balance is achieved with fewer noise and richer spatial information, effectively preserving spatial coherence.